OpenStudy (anonymous):

what is the solution for x+3y=6 6x-9y=0

OpenStudy (xguardians):

What are you solving for? X or y?

OpenStudy (anonymous):

system of equations

OpenStudy (luigi0210):

You could solve it by graphing, substitution or solving by eliminating one variable.

OpenStudy (anonymous):

what is the solution for the system of equations for x+3y=6 and 6x-9y=0

OpenStudy (triciaal):

using substitution first eqn x = 6-3y use this expression for x in 2nd eqn 6(6-3y)-9y = 0 36-18y-9y = 0 36 = 27y y =4/3 x = 6 - 3*4/3 =2 check values in eqn 2 6x -9y = 0 6*2 - 9 *4/3 = 0

OpenStudy (triciaal):

using elimination multiply the first eqn by 6 6x -18y =36 subtract the 2nd eqn 6x -6x + 18y -(-9y) =36 - 0 27y = 36 divide by 27 y = 4/3 from eqn 1 putting in your y value x + 3 (4/3) = 6 x = 6 - 4 = 2